OpenStudy (idealist10):

Solve the initial-value problem y'+(2/x)y=-2xy/(x^2+2x^2*y+1), y(1)=-2.

OpenStudy (idealist10):

@hartnn @amistre64

OpenStudy (anonymous):

\[\begin{align*} y'+\frac{2}{x}y&=-\frac{2xy}{x^2+2x^2y+1}\\\\ y'&=-\frac{4x^2y+4x^2y^2+2y}{x(x^2+2x^2y+1)}\\\\ 4x^2y+4x^2y^2+2y+x(x^2+2x^2y+1)y'&=0 \end{align*}\] Check for exactness: \[\frac{\partial }{\partial y}\left[4x^2y+4x^2y^2+2y\right]=4x^2+8x^2y+2\\ \frac{\partial }{\partial x}\left[x^3+2x^3y+x\right]=3x^2+6x^2y+1\] Integrating factor: \[\begin{align*}\mu(x)&=\exp\left(\int\frac{(4x^2+8x^2y+2)-(3x^2+6x^2y+1)}{x^3+2x^3y+x}~dx\right)\\\\ &=\exp\left(\int\frac{x^2+2x^2y+1}{x^3+2x^3y+x}~dx\right)\\\\ &=\exp\left(\int\frac{dx}{x}\right)\\\\ &=x \end{align*}\] Distribute the IF: \[\begin{align*} 4x^3y+4x^3y^2+2xy+(x^4+2x^4y+x^2)y'&=0 \end{align*}\] Check for exactness again: \[\frac{\partial }{\partial y}\left[4x^3y+4x^3y^2+2xy\right]=4x^3+8x^3y+2x\\ \frac{\partial }{\partial x}\left[x^4+2x^4y+x^2\right]=4x^3+8x^3y+2x\] Now you're ready to solve for \(\Psi(x,y)\). \[\begin{align*}\Psi_x&=4x^3y+4x^3y^2+2xy\\\\ \Psi&=\int\cdots~dx\\\\ &=x^4y+x^4y^2+x^2y+f(y)\\\\ \Psi_y&=x^4+2x^4y+x^2+f'(y)\\\\ x^4+2x^4y+x^2&=\\\\ f'(y)&=0\\\\ f(y)=C\end{align*}\]

OpenStudy (idealist10):

Thank you!

OpenStudy (anonymous):

You're welcome! I'm assuming the exact approach is fine by you? I was looking for a proper substitution the first time around, but I saw you posting questions about exact equations lately so I figured you wouldn't have trouble following this one.